OpenStudy (anonymous):

the price of a delta airline ticket from los angeles to boston increased to $400. This is a 15% increase. what was the old fare?

OpenStudy (shamim):

100-15=?

OpenStudy (shamim):

85/100=0.85?

OpenStudy (shamim):

Now 0.85*400=?

OpenStudy (amorfide):

when using percentages to increase a number by 15 percent to equal 400 you have \[\frac{ 15 \times x }{ 100 }+ x=400\]

OpenStudy (amorfide):

rearrange to solve for x

OpenStudy (amorfide):

where x is the starting price, 15/100 is 15% of the original number and 400 is the price it currently is you could also do it another way knowing it was increased by 15% then you know that 1 is 100% so 1+0.15 is 115% meaning x+15% of x so 1.15x=400 solve for x

OpenStudy (amorfide):

hope that makes it clearer for you :) btw Shamim is wrong, as it would only reduce your current number by 15%, it wouldn't give you the original number that was increased by 15%

OpenStudy (amorfide):

@u009574

OpenStudy (amorfide):

please tell me if you need help still or something explained